# Hyperspectral Image Generation with Unmixing Guided Diffusion Model

### v1

@@ V Conclusion and Discussion
This paper introduces an unconditional hyperspectral image generation model that achieves both high fidelity and diversity, addressing the high dimensionality and physical constraint challenges posed by hyperspectral data. To overcome these challenges, we propose a novel framework comprising a hyperspectral unmixing autoencoder pair, which projects the original HSI into a low-dimensional abundance space for efficient generation. Additionally, we design a diffusion module capable of generating high-quality samples while satisfying the physical constraints of non-negativity and unity. Furthermore, we introduce a new evaluation method specifically tailored to the characteristics of hyperspectral data, providing a more accurate assessment of generation quality. Experimental results demonstrate that our model outperforms existing methods, generating higher-quality HSIs and paving the way for future advancements in hyperspectral image generation research.
HUD makes full use of hyperspectral unmixing to guide the latent space of the diffusion model. However, constrained by the characteristics of HSI that have limited data and significant distribution differences, we do not recommend further training the unmixing-initialized autoencoder to avoid severe overfitting. Additionally, we can only train one model per image rather than generating different hyperspectral images from a unified diffusion model. Furthermore, despite the significant reduction in dimensionality achieved by unmixing guidance, it is still not feasible to generate complete high-resolution hyperspectral images in practice. To address these issues, our future work will focus on a different methodology that can generate images with different bands. Additionally, spatial compression is also under consideration.

### v3

@@ V Conclusion and Discussion
We presented an unconditional hyperspectral image (HSI) generation framework that jointly advances fidelity and diversity while addressing two core challenges of HSI synthesis: high spectral dimensionality and physics-driven constraints. Our approach couples a hyperspectral unmixing autoencoder with a diffusion-based generator. The unmixing module projects HSIs into a low-dimensional abundance space, enabling efficient modeling, and the diffusion module operates in a constraint-relaxed domain with a differentiable projection that guarantees non-negativity and sum-to-one upon reconstruction. We further introduced HSI-specific evaluation metrics that better reflect the spectral-spatial characteristics of unconditional generation. Empirically, our model surpasses existing approaches across both proposed and conventional metrics, indicating its effectiveness for realistic and diverse HSI synthesis.
HUD leverages unmixing to impose a physically meaningful latent structure for diffusion. Nevertheless, the data-sparse and distribution-shifted nature of HSI collections poses practical limitations. To mitigate overfitting, we refrain from extensively fine-tuning the unmixing-initialized autoencoder, which constrains adaptability across scenes. Moreover, under current data and computational limit, we train models on a per-image (or per-scene) basis rather than a single, universal generator. The full-resolution, large-footprint HSI synthesis remains computationally prohibitive despite dimensionality reductions. Future work will pursue: (1) A unified generative paradigm capable of accommodating varying numbers of spectral bands and sensor characteristics. (2) Enhanced spatial modeling, including principled spatial compression and multiscale tiling strategies to approach full-resolution synthesis. (3) Data-efficient training protocols (e.g., cross-scene pretraining, self-supervision, and physics-informed regularization) to broaden generalization while preserving physical consistency.